OpenStudy (anonymous):

Given the function f(x) = x2 and k = –3, which of the following represents a vertical shift?

OpenStudy (anonymous):

f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (solomonzelman):

when you are dealing with any function \(f(x)=x^w\), a vertical shift is: \(f(x)=x^w+c\) (shift up), OR \(f(x)+x^w-c\) (shift down).

OpenStudy (solomonzelman):

does this help, or not?

OpenStudy (solomonzelman):

I mean \(fx)=x^w-c\) for the last equation ...

OpenStudy (solomonzelman):

btw, you don't need to know the value of the k to answer this question.

OpenStudy (anonymous):

yes this helps.

OpenStudy (solomonzelman):

Ok, and your answer is?

OpenStudy (anonymous):

would it be d?

OpenStudy (solomonzelman):

I guess it didn't help at all....

OpenStudy (solomonzelman):

it seems as though you are guessing....

OpenStudy (anonymous):

I'm trying to understand

OpenStudy (anonymous):

okay wait would it be a?

OpenStudy (anonymous):

im so bad at math sorry I'm really trying to get this

OpenStudy (solomonzelman):

Say you have a function \(f(x)=x^w\) -------------------------------- \(f(x\cdot k)=(x\cdot k)^w\) that is just increasing the base, and gives you a different function. \(f(x)+c=x^w+c\) ~ shift up when c>0 ~ shift down when c<0. \(kf(x)=k(x)^w\) ~ you are stretching the function when k>1, ~ shrinking when 0<k<1, ~ reflecting the function across x axis and then shrinking it when -1<k<0, ~ reflecting the function across x axis and stretching it when k<-1. \(f(x+k)=(x+k)^w\) ~ shift right when k<0 ~ shift left when k>0

OpenStudy (anonymous):

so it would be a

OpenStudy (solomonzelman):

yes, A is the correct answer.

OpenStudy (anonymous):

thank you! it was right.

OpenStudy (solomonzelman):

:)